# IMUFace: Real-Time, Low-Power, Continuous 3D Facial Reconstruction Through Earphones

### v1

@@ III-B Sensing Principle
Figure 3: Device Prototype: Composed of a pair of earbuds and XLAO nRF52840 Sense MCU.
IV Method
In Fig. 2, we present an overview of our proposed IMUFace system. First, we capture changes in facial expressions using inertial measurement units (IMUs) embedded in earphones, in conjunction with corresponding video data of the human face. The collected data undergoes preprocessing. A quick calibration process is conducted for the IMU data to ensure that the IMUs effectively capture skin movement. Next, we apply the Short-Time Fourier Transform (STFT) to extract frequency domain features from the IMU data, after which we segment this data.
To process the video data, we begin by extracting facial landmarks and normalizing them. These normalized landmarks serve as the ground truth during the training process. Next, we input the preprocessed IMU data into a transformer-based deep learning model. This model uses information from the IMU sensors and analyzes the relationships between consecutive frames to estimate the positions of facial landmarks in the current frame. Finally, we fit these landmarks to the parameters of the FLAME model (Faces Learned with an Articulated Model and Expressions) to create 3D facial reconstruction animations.
IV-A Hardware Prototype Design
@@
IV-B Data Processing Pipeline
IV-B1 Calibration
@@
IV-B2 Signal Synchronization and Filtering
@@
IV-B3 Signal Normalization
@@
First, we select the nose tip as the origin and shift all landmarks accordingly:
@@
where $P_{i}=(x_{i},y_{i})$ are the original landmark coordinates, and $(x_{\text{nose}},y_{\text{nose}})$ are the coordinates of the nose tip.
Next, by rotating the coordinates, we align the $x$-axis to the direction from the left outer eye corner to the right outer eye corner. The rotation angle $\theta$ is calculated as:
@@
where $(x_{\text{left}},y_{\text{left}})$ and $(x_{\text{right}},y_{\text{right}})$ represent the coordinates of the left and right outer eye corners, respectively.
The rotated coordinates $P_{i}^{\prime\prime}$ are obtained using the following transformation:
@@
Finally, we normalize the distance between the outer eye corners to 1, ensuring uniform scaling. The normalization factor $d$ is given by:
@@
The final normalized coordinates $P_{i}^{\prime\prime\prime}$ are computed as:
@@
Figure 4: Illustration of facial landmark alignment.
IV-B4 Short-Time Fourier Transform
Each facial expression movement is associated with unique frequency components. Since the STFT better isolates noise components and enhances the clarity of extracted features from time-series data [38], we use STFT to convert time-domain signals into time-frequency domain spectrograms. The STFT of a signal $x(t)$ is computed as:
[…]
where $w(\tau-t)$ is a window function centered at $t$, $f$ is the frequency, and $X(t,f)$ represents the time-frequency representation of the signal. Frequency-domain features improve the training of our model. We apply STFT to each channel of the IMU data with a window length of 30, a step size of 1, and 32 frequency components. Considering the model complexity, we concatenate six spectrograms (sharing the same time axis) to form a single feature representation input. The concatenation is along the frequency axis, resulting in a final frequency-domain feature dimension of $12\times 17=204$, with a length of IMU signal length.
IV-B5 Data Segmentation
We collect IMU signal data with 12 channels and extract frequency-domain features for each channel. Both the time-domain data and the extracted frequency-domain features are used as inputs to the model. To enable the model to utilize rich temporal relationships when predicting facial landmarks, we input ten continuous frames of data, each with a size of 1 × 216.
@@ IV-B5 Data Segmentation
IV-C Model Architecture and Training Strategy
In this study, we designed a deep-learning model called IMUTwinTrans. This model, based on a ConvTransformer architecture, is capable of simultaneously leveraging the temporal and frequency features of IMU signals to establish the correspondence between IMU signals and facial landmarks, as illustrated in Fig. 5.
IV-C1 CNN-based Feature Extraction
@@
IV-C2 Transformer Structure
We perform value embedding and position encoding on the feature data derived from the CNN output (size 10×512) and feed it into a two-layer Transformer Encoder. The encoder includes a self-attention module within the multi-head attention layers, a multilayer perceptron, and normalization layers. Residual connections are incorporated at both the multi-head attention and multilayer perceptron layers to boost the convergence of the machine learning model. Regarding model parameters, we set $d_{model}$ to 512, $n_{head}$ to 4, $d_{ff}$ to 1024, and the dropout ratio to 0.1 in the encoder.
IV-C3 Loss Function
In the task of facial landmark prediction, standard MSE and MAE loss functions treat all errors equally, regardless of their magnitude. In particular, MSE heavily penalizes large errors while neglecting small ones, resulting in important but less active landmarks (such as pupils compared to lips) not receiving sufficient attention during training. To address this issue, we adopt the Wing Loss function [39], which applies a relatively smooth treatment to small errors while assigning higher weights to large errors. This helps the model predict points closer to the true values more accurately without overly penalizing the model for a few large errors. The loss for each facial landmark is defined as:
Wing $\text{Wing Loss}(x_{i})=\begin{cases}w\log(1+\frac{|x_{i}|}{\epsilon}),&\text{if}|x_{i}|<w\\ |x_{i}|-C,&\text{if}|x_{i}|\geq w\end{cases}$ (7)
@@ IV-C4 Optimization
IV-D 3D Reconstruction
To improve the system’s usability and reduce the modeling’s complexity, we seek a compact head model that can easily fit the data while preserving enough details to generate expressive facial animations.
Figure 6: Performance of continuous facial landmark tracking for each participant.
FLAME 3D Head Model. IMUFace uses the FLAME [40] model to represent the face in 3D. FLAME is a statistical head model that learns from approximately 33,000 3D scans. It combines a PCA space for identity shape, simple rotational degrees of freedom for the neck, jaw, and eyeballs, pose-dependent corrective blend shapes, and global expression blend shapes to efficiently represent complex facial geometry and expressions. Mathematically, the model is defined as:
@@
Optimization-based Shape Reconstruction. We employ a series of optimization steps to obtain a 3D face model that closely resembles the user’s real expressions. First, we perform camera calibration by optimizing the parameters for scale, rotation, and translation to minimize the L2 distance between the 2D projections of the 3D face model’s facial landmarks and the user’s facial landmarks. Second, we optimize the FLAME model parameters (shape, pose, and expression) to further reduce the L2 distance between the 2D projections of the facial landmarks from the 3D face model and the user’s facial landmarks. Ultimately, we can render the user’s 3D face model using the optimal FLAME model parameters.
V Performance Evaluation
@@
V-A Experimental Methodology
V-A1 Data Collection
In the study, we recruited 12 participants to evaluate IMUFace’s performance. The participants included nine males and three females from the local institution, aged from 18 to 28. The experimental equipment comprises the IMUFace prototype, a Hikvision camera, and a desktop computer. Participants were asked to wear the IMUFace wireless earphones prototype, which was connected to the computer via Bluetooth.
@@ V-A1 Data Collection
Each participant completed 12 sessions of the experiment. In each session, the participants were required to make seven expressions, which were prompted in a random order. Each expression was made twice. Each data collection session lasted for 60 seconds, and the participants took adequate rest between sessions. The earphone prototype was removed and remounted across sessions to validate the robustness to natural changes in sensor positions during daily usage.
V-A2 Experimental Settings
@@
V-A3 Evaluation Metric
The IMUFace system is evaluated primarily by two metrics [19]:
1) Mean Absolute Error (MAE): the absolute difference between the predicted facial landmarks and the ground-truth facial landmarks. It is defined as: $MAE=\frac{1}{N}\times\sum\|g-r\|\times\frac{d_{real}}{d_{norm}}$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. ${d_{real}}$ denotes the actual outer corner distance in millimeters(we measured ${d_{real}}$ for each user who participated in the study). ${d_{norm}}$ denotes the outer corner distance in the normalized coordinate frame. From Section IV-B3, we know that ${d_{norm}}=1$. This provided the MAE in units of millimeters, which is the primary metric used to evaluate IMUFace.
2) Normalized Mean Error (NME): the average error between the predicted facial landmarks and the ground truth facial landmarks, represented by $NME=\frac{1}{N}\times\sum\frac{\|g-r\|}{d_{norm}}\times 100\%$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. ${d_{norm}}$ denotes the outer corner distance in the normalized coordinate frame. From Section IV-B3, we know that ${d_{norm}}=1$.
@@ V-A3 Evaluation Metric
TABLE I: Comparison result of different Models.
@@
Bi-LSTM [41] 2.57 mm 3.97%
TCN [42] 2.39 mm 3.68%
IF-ConvTransformer [43] 2.27 mm 3.50%
Mamba2 [44] 2.37 mm 3.64%
@@
V-B User-Dependent Evaluation
For each user’s 12 sessions data, we randomly selected five sessions for training. The rest was used for testing in a randomized cross-validation fashion. Fig. 6(a) illustrates the average MAE & NME and corresponding standard deviations for all the 51 facial landmarks of each participant. Specifically, IMUFace obtains an average of 2.21 mm MAE and 3.40% NME with average standard deviations of 0.99 mm and 0.90%. Among all participants, U4 performed the best, with an MAE of only 1.46, while U8 had the highest MAE at 2.72. Fig. 6(b) depicts the Cumulative Density Function (CDF) of the MAE errors for each individual participant as well as cross-participant cases. 60.0% of the MAEs are less than 2.21, indicating a high level of prediction accuracy. Moreover, different landmarks may exhibit varying error levels due to their differing movement characteristics. Fig. 6(c) visualizes the average MAE for all 51 major landmarks. We found that the reconstructed landmarks on the eyebrows and mouth have relatively more significant errors, while the facial landmarks around the eyes have more minor errors. This is mainly because the eyebrows and mouth change more during facial expression variations.
V-C User-Adaptive Evaluation
V-D Micro-benchmark Tests
V-D1 Model Selection
To verify the effectiveness of our model, we compared it with several deep learning network methods commonly used for IMU data analysis to demonstrate the superiority of our model. We evaluated the performance of different networks under user-dependent experimental settings, and the results are presented in Table I.
@@ V-D2 Number of Earphones
V-D3 Impact of Training Data Size
@@
V-D4 Latency and Power Consumption

### v3

@@ 4 Method
Figure 3: Device Prototype: Composed of a pair of earbuds and XIAO nRF52840 Sense MCU.
4.1 Hardware Prototype Design
@@
4.2 Data Processing Pipeline
4.2.1 Calibration
@@
4.2.2 Signal Synchronization and Filtering
@@
Figure 4: Illustration of facial landmark alignment. (a) Before landmark alignment. (b) After landmark alignment.
4.2.3 Signal Normalization
@@
First, we select the nose tip as the origin and shift all landmarks accordingly, as defined in Eq. (1):
@@
where $P_{i}=(x_{i},y_{i})$ denotes the original landmark coordinates, and $(x_{\text{nose}},y_{\text{nose}})$ represents the position of the nose tip.
Next, we perform a coordinate rotation to align the $x$-axis with the direction from the left outer eye corner to the right outer eye corner. The rotation angle $\theta$ is computed using Eq. (2):
@@
where $(x_{\text{left}},y_{\text{left}})$ and $(x_{\text{right}},y_{\text{right}})$ correspond to the coordinates of the outer corners of the left and right eyes, respectively.
The rotated landmark coordinates $P_{i}^{\prime\prime}$ are then obtained by applying the transformation in Eq. (3):
@@
To ensure uniform scaling, we normalize the distance between the two outer eye corners to 1. The normalization factor $d$ is calculated according to Eq. (4):
@@
Finally, the normalized landmark coordinates $P_{i}^{\prime\prime\prime}$ are obtained by applying Eq. (5):
@@
4.2.4 Short-Time Fourier Transform
Each facial expression movement is associated with unique frequency components. Since the STFT better isolates noise components and enhances the clarity of extracted features from time-series data [zhang2022force], we use STFT to convert time-domain signals into time-frequency domain spectrograms. The STFT of a signal $x(t)$ is computed as Eq. (6):
$X(t,f)=\int_{-\infty}^{\infty}x(\tau)w(\tau-t)e^{-j2\pi f\tau}\,d\tau$ (6)
where $w(\tau-t)$ is a window function centered at $t$, $f$ is the frequency, and $X(t,f)$ represents the time-frequency representation of the signal. Frequency-domain features improve the training of our model. We apply STFT to each channel of the IMU data with a window length of 30, a step size of 1, and 32 frequency components. Considering the model complexity, we concatenate six spectrograms (sharing the same time axis) to form a single feature representation input. The concatenation is along the frequency axis, resulting in a final frequency-domain feature dimension of $12\times 17=204$, with a length of IMU signal.
@@ 4.2.4 Short-Time Fourier Transform
4.2.5 Data Segmentation
We collect IMU signal data with 12 channels and extract frequency-domain features for each channel. Both the time-domain data and the extracted frequency-domain features are used as inputs to the model. To help the model capture temporal relationships when predicting facial landmarks, we use ten consecutive frames of processed IMU data as one input segment.
4.3 Model Architecture and Training Strategy
In this study, we designed a deep-learning model called IMUTwinTrans. This model, based on a ConvTransformer architecture, leverages both temporal and frequency features of IMU signals to establish the correspondence between IMU data and facial landmarks, as illustrated in Fig. 5. The model input consists of 10 consecutive frames of IMU data that include both time-domain and frequency-domain features. Specifically, the time-domain signal for each frame is arranged as a $30\times 12$ matrix. The frequency-domain features are obtained via a Short-Time Fourier Transform (STFT), resulting in a $30\times 204$ matrix. By concatenating the time-domain and frequency-domain features, each frame forms a combined feature matrix of $30\times 216$. The input data are first processed by a Convolutional Neural Network (CNN) for feature extraction, followed by a Transformer encoder to capture temporal dependencies. Finally, a linear projection layer outputs the predicted coordinates of 51 facial landmarks, represented as a $1\times 102$ vector.
4.3.1 CNN-based Feature Extraction
@@
4.3.2 Transformer Structure
We perform value embedding and position encoding on the feature data derived from the CNN output (size $10\times 512$) and feed it into a two-layer Transformer Encoder. The encoder includes a self-attention module within the multi-head attention layers, a multilayer perceptron, and normalization layers. Residual connections are incorporated at both the multi-head attention and multilayer perceptron layers to boost the convergence of the machine learning model. Regarding model parameters, we set $d_{model}$ to 512, $n_{head}$ to 4, $d_{ff}$ to 1024, and the dropout ratio to 0.1 in the encoder.
4.3.3 Loss Function
In the task of facial landmark prediction, standard MSE and MAE loss functions treat all errors equally, regardless of their magnitude. In particular, MSE heavily penalizes large errors while neglecting small ones, resulting in important but less active landmarks (such as pupils compared to lips) not receiving sufficient attention during training. To address this issue, we adopt the Wing Loss function [feng2018wing], which applies a relatively smooth treatment to small errors while assigning higher weights to large errors. This helps the model predict points closer to the true values more accurately without overly penalizing the model for a few large errors. The loss for each facial landmark is defined as (Eq. (7)):
$\text{Wing Loss}(x_{i})=\begin{cases}w\log(1+\frac{|x_{i}|}{\epsilon}),&\text{if}|x_{i}|<w\\ |x_{i}|-C,&\text{if}|x_{i}|\geq w\end{cases}$ (7)
@@ 4.3.4 Optimization
4.4 3D Reconstruction
To enhance system usability and simplify the modeling process, we aim to use a compact head model that balances reduced complexity with sufficient detail to produce expressive facial animations.
The IMUFace system adopts the FLAME model (2020 version) for 3D facial representation. Specifically, the 3D reconstruction is primarily carried out using the FLAME PyTorch framework. FLAME is a statistical head model built from a dataset of approximately 33,000 3D scans. It efficiently captures complex facial geometries and expressions through a combination of identity PCA space, rotational degrees of freedom for the neck, jaw, and eyes, pose-dependent corrective blend shapes, and global expression blend shapes. The model is mathematically represented as (Eq. (8)):
@@
A series of optimization steps is implemented to reconstruct a 3D face model that accurately reflects the user’s expressions. The process begins with camera calibration, where the parameters for scale, rotation, and translation are optimized to minimize the L2 distance between the 2D projections of the 3D model’s facial landmarks and the user’s facial landmarks. Following this, the FLAME model parameters—covering shape ($\vec{\beta}$), pose ($\vec{\theta}$), and expression ($\vec{\psi}$)—are fine-tuned to further reduce the L2 distance between the projected landmarks of the 3D model and the user’s facial landmarks. This L2-distance-based optimization method enables the model to fully adapt and reconstruct a unique shape that closely matches the target user’s facial features.
Using the optimized FLAME parameters, the user’s 3D face model can then be rendered with high accuracy. While we customize the shape parameters for each user to capture their unique head geometry, we uniformly apply FLAME’s default texture parameters. This decision was primarily driven by our application’s focus on the dynamic representation of facial expressions and the accuracy of geometric shapes. The default texture provides a sufficient basic visual effect while avoiding the additional complexity and data acquisition costs associated with obtaining and fitting individual textures. Furthermore, a unified texture simplifies the rendering process, contributing to the system’s consistency and robustness.
5 Performance Evaluation
@@
5.1 Experimental Methodology
5.1.1 Data Collection
In the study, we recruited 12 participants to evaluate IMUFace’s performance. The participants included nine males and three females from the local institution, aged 18 to 28. The experimental equipment comprises the IMUFace prototype, a Hikvision camera, and a desktop computer. Participants were asked to wear the prototype of the IMUFace wireless earphones, which were connected to the computer via Bluetooth. This study was reviewed and approved by the Ethics Review Committee (ERC) at the South China University of Technology.
@@ 5.1.1 Data Collection
Each participant completed 12 sessions of the experiment. In each session, the participants were required to make seven expressions, which were prompted in a random order. Each expression was made twice. Each data collection session lasted 60 seconds, and the participants took adequate rest between sessions. The earphone prototype was removed and remounted across sessions to validate the robustness to natural changes in sensor positions during daily usage.
5.1.2 Experimental Settings
@@
5.1.3 Evaluation Metric
The IMUFace system is evaluated primarily by two metrics [wu2021bioface]:
1) Mean Absolute Error (MAE): the absolute difference between the predicted facial landmarks and the ground-truth facial landmarks. It is defined as: $MAE=\frac{1}{N}\times\sum\|g-r\|\times\frac{d_{real}}{d_{norm}}$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. $d_{real}$} denotes the actual outer corner distance in millimeters(we measured $d_{real}$} for each user who participated in the study). $d_{norm}$} denotes the outer corner distance in the normalized coordinate frame. From Section 4.2.3, we know that ${d_{norm}}=1$. This provided the MAE in units of millimeters, which is the primary metric used to evaluate IMUFace.
2) Normalized Mean Error (NME): the average error between the predicted facial landmarks and the ground truth facial landmarks, represented by $NME=\frac{1}{N}\times\sum\frac{\|g-r\|}{d_{norm}}\times 100\%$, where $g$ and $r$ denote the facial landmarks extracted from the ground truth of vision and IMUFace in a normalized coordinate frame. $d_{norm}$} denotes the outer corner distance in the normalized coordinate frame. From Section 4.2.3, we know that ${d_{norm}}=1$.
Figure 6: Performance of continuous facial landmark tracking for each participant. (a) Per-participant Landmark Tracking Error. (b) MAE CDF for each participant. (c) Visualization of average MAE.
5.2 User-Dependent Evaluation
For each user’s 12 recorded sessions, we randomly selected five sessions for training, while the remaining sessions were used for testing in a randomized cross-validation manner. Fig. 6(a) presents the average Mean Absolute Error (MAE) and Normalized Mean Error (NME), along with their corresponding standard deviations, for all 51 facial landmarks of each participant. Specifically, the blue bars represent the average MAE (in millimeters), which reflects the absolute spatial deviation between the predicted and ground truth landmarks. The red bars represent the average NME (in percentage), which measures the relative deviation normalized by the facial size, enabling cross-user performance comparison regardless of individual facial dimensions. These two complementary metrics provide a comprehensive evaluation of both absolute and relative reconstruction accuracy. IMUFace achieves an average MAE of 2.21 mm and an NME of 3.40%, with standard deviations of 0.99 mm and 0.90%, respectively. Among all participants, U4 achieves the best performance with an MAE of only 1.46 mm, while U8 exhibits the highest MAE at 2.72 mm. Fig. 6(b) illustrates the Cumulative Density Function (CDF) of MAE for each individual participant as well as across all participants. The CDF curve reveals that 60.0% of the samples have an MAE below 2.21 mm, indicating a high overall prediction accuracy. This distribution further confirms the robustness of our model in reconstructing facial landmarks with low error rates across different users. In addition, different landmarks tend to exhibit varying error levels due to their distinct movement patterns during facial expressions. Fig. 6(c) visualizes the average MAE for each of the 51 facial landmarks. We observe that landmarks located on the eyebrows and mouth tend to have higher reconstruction errors, while those around the eyes generally exhibit lower errors. This is mainly because the eyebrows and mouth undergo more dynamic deformations during expression changes, making their accurate reconstruction more challenging.
5.3 User-Adaptive Evaluation
@@
Table 1: Comparison result of different Models.
@@
Bi-LSTM [zhang2015bidirectional] 2.57 mm 3.97%
TCN [verma2021expressear] 2.39 mm 3.68%
IF-ConvTransformer [zhang2022if] 2.27 mm 3.50%
Mamba2 [dao2024transformers] 2.37 mm 3.64%
@@
5.4 Micro-benchmark Tests
5.4.1 Model Selection
To verify the effectiveness of our model, we compared it with several deep learning network methods commonly used for IMU data analysis to demonstrate the superiority of our model. We evaluated the performance of different networks under user-dependent experimental settings, and the results are presented in Table 1.
@@ 5.4.2 Number of Earphones
5.4.3 Impact of Training Data Size
@@
While the average MAE between two and five minutes of training data appears numerically close (2.27 mm vs. 2.21 mm), a more detailed analysis reveals a noticeable improvement in the normalized mean error (NME), which decreases from 3.65% to 3.40%. Since NME accounts for differences in face scale and user variation, this reduction suggests that the model trained with five minutes of data has better generalization across diverse users and face shapes. This improvement in normalized prediction accuracy justifies the use of a longer training duration. Additionally, the extended data collection enhances the model’s robustness and stability in handling complex facial dynamics, making it more reliable for real-world applications.
5.4.4 Comparison with Existing Solutions
To compare the performance of IMUFace with existing vision-based and wearable methods, Table 2 presents the NME results across a range of representative systems. SDM [zhu2015face], CFSS [xiong2013supervised], and HRNet [wang2020deep] are classic vision-based approaches that rely on public image datasets such as 300-W to track 68 facial landmarks. In contrast, mm3DFace [xie2023mm3dface] and BioFace-3D [wu2021bioface] adopt alternative sensing modalities—millimeter-wave and EMG signals respectively—on self-collected datasets. EARFace [zhang2023earphone] is another non-visual method that uses in-ear sensing by emitting and receiving acoustic signals in the ear canal.
Our method, IMUFace, achieves an NME of 3.40% on a self-collected dataset with 51 landmarks, which is very close to the performance of BioFace-3D (3.38%). However, unlike BioFace-3D, which requires attaching electrodes directly to the user’s face, IMUFace relies only on lightweight wearable IMU sensors, making it significantly more comfortable and user-friendly.
Although IMUFace’s accuracy is slightly lower than EARFace (3.14%), it avoids several practical limitations of the EARFace system. Specifically, EARFace relies on wired earphones and uses active acoustic signal emission into the ear canal, which may pose risks to user hearing health, involves higher power consumption, and compromises convenience due to its tethered design. In contrast, IMUFace offers comparable performance while being low-power, wireless, and non-invasive. These advantages highlight its strong potential for practical deployment, particularly in scenarios where visual input is unavailable or privacy-sensitive environments are involved.
Table 2: Comparison with existing solutions.
Methods Dataset Landmarks NME (%)
SDM [zhu2015face] 300-W 68 7.52
CFSS [xiong2013supervised] 300-W 68 5.76
HRNet [wang2020deep] 300-W 68 2.87
mm3DFace [xie2023mm3dface] Self-collected 68 3.94
BioFace-3D [wu2021bioface] Self-collected 53 3.38
EARFace [zhang2023earphone] Self-collected 51 3.14
IMUFace (Ours) Self-collected 51 3.40
5.4.5 Latency and Power Consumption